Assume the random variable X is normally distributed with mean mu equals 50 and standard deviation sigma equals 7. Compute the probability. Be sure to draw a normal curve with the area corresponding to the probability shaded. P(X>38)=
Answer #1

Solution Given! XNNcu6²) u=50 6= 7 p(x>38) = - p(x533) = 1- P1 xy < 38-50) =l-P(Z -1.7143 ) from table =) - 0.0432 = 0.9568 p